# Plasma Mixing in Collisionless Kelvin–Helmholtz Dynamics

### v1

@@ Abstract
Aims. We investigate the efficiency and physical mechanisms of plasma mixing driven by the nonlinear evolution of the KHI.
Methods. We perform high-resolution two-dimensional Particle-In-Cell (PIC) simulations using a finite-Larmor-radius shear-flow initial configuration. Plasma mixing is quantified using particle tracking, passive tracers, and diagnostics of magnetic reconnection.
Results. Mixing across the shear layer is present but localized, occurring mainly in narrow interface regions and plasma structures. Ions mix more effectively than electrons, which remain largely frozen to field lines. Enhanced mixing correlates with localized reconnection within and between KH vortices.
Conclusions. Cross-boundary transport driven by the kinetic KHI is highly localized and mediated by vortex advection and reconnection. Electron mixing is strongly constrained, providing an upper bound on kinetic-scale transport across collisionless shear layers.
@@ 1 Introduction
The Kelvin–Helmholtz instability (KHI) develops when a plasma shear flow becomes super-Alfvénic with respect to the magnetic-field component parallel to the shear (Chandrasekhar 1961). At Earth’s low-latitude magnetopause, this condition is frequently met during intervals of strongly northward or southward interplanetary magnetic field (IMF), and signatures of surface waves and rolled-up vortices have been widely observed (Hasegawa et al. 2004; Eriksson et al. 2016; Stawarz et al. 2016; Settino et al. 2024). KH activity has also been reported at other planetary magnetospheres, including Mercury (Aizawa et al. 2020), Mars (Koh et al. 2025), Saturn (Dialynas 2018), and Jupiter (Montgomery et al. 2023), raising the question of how magnetospheric size influences the multi-scale development of the instability. After saturation, KH vortices generate strong gradients in velocity, density, and magnetic field down to ion and electron scales, triggering secondary instabilities that broaden the mixing layer and enhance plasma transport across the solar wind–magnetosphere boundary (Nykyri and Otto 2001; Hasegawa et al. 2004; Nakamura et al. 2017a). Kinetic studies and in-situ observations show that this transport is mediated by processes such as reconnection, wave–particle interactions, and vortex-induced reconnection (VIR), enabling solar-wind plasma entry even under northward IMF conditions (Nykyri et al. 2006; Stawarz et al. 2016; Nakamura et al. 2017b, 2020). While magnetohydrodynamic (MHD) simulations capture the global KH dynamics and reconnection-driven transport (Faganello et al. 2012; Ma et al. 2017; Ferro et al. 2024), they do not resolve the kinetic mechanisms governing particle diffusion and heating. Despite these advances, quantitatively characterizing plasma mixing in fully kinetic KHI remains challenging, particularly in distinguishing genuine plasma interpenetration from large-scale advection (Nakamura et al. 2013; Karimabadi et al. 2013).
In this Letter, we investigate plasma mixing driven by the nonlinear evolution of the collisionless KHI using fully kinetic Particle-in-Cell (PIC) simulations of an electron–ion plasma. We quantify mixing efficiency based on particle origin and kinetic diagnostics, and discuss the implications for solar wind–magnetosphere interactions.
@@ 2 Methodology and Simulation Setup
We model a collisionless magnetized plasma with uniform per-species density $n_{0}$ and an initial magnetic field composed of an in-plane component along $x$ and a guide field $B_{z,0}=10B_{x,0}$. All quantities are normalized using ion reference scales: the ion gyro-frequency $\Omega_{c,i}=q_{i}B_{\mathrm{tot,0}}/(m_{i}c)$, where $B_{\mathrm{tot,0}}=\sqrt{B_{x,0}^{2}+B_{z,0}^{2}}$ is the total magnetic field magnitude at initialization; the ion plasma frequency $\omega_{p,i}=(4\pi n_{0}q_{i}^{2}/m_{i})^{1/2}$; the ion inertial length $d_{i}=c/\omega_{p,i}$; and the Alfvén speed $v_{A}=B_{\mathrm{tot,0}}/\sqrt{4\pi n_{0}m_{i}}$. At initialization, the ion cyclotron-to-plasma frequency ratio is $\Omega_{c,i}/\omega_{p,i}=0.05$.
@@
3 Plasma Mixing and Magnetic Reconnection
In Figure 1, we show snapshots representative of the evolution of the vortices at the lower shear during the nonlinear stage of the KHI. The quantity shown is the electron mixing fraction (Nakamura and Daughton 2014), $F_{e}=(n_{e,\ell=1}-n_{e,\ell=0})/n_{e}$, where $n_{e,\ell=1}$ and $n_{e,\ell=0}$ are the electron number densities initially on either side of the shear. Regions with $F_{e}=-1$, $F_{e}=1$, and $|F_{e}|<1$ correspond to unmixed plasma on either side and mixed plasma, respectively. The ion mixing fraction and the upper shear layer for both species behave similarly and are not shown.
The left panel of Figure 1 shows three rolled-up vortices at the early nonlinear stage. As they merge (middle panel), a dynamic mixing layer forms, where plasma parcels can temporarily cross regions before returning to their origin (Figure 1, middle and left panels). Magnetic tension limits permanent cross-field transport: although reconnection allows plasma parcels to detach from their original field lines, the strong in-plane field pulls them back. Consequently, magnetic tension counteracts plasma transport and forces macroscopic parcels to rebound toward their region of origin, limiting their ability to permanently migrate into the opposite region. Appendix B further illustrates how the in-plane field constrains KHI-driven mixing.
While the mixing fraction provides the qualitative picture described above, it does not allow us to adequately quantify actual mixing because it does not capture the movement of small groups of particles potentially crossing between regions. To quantify interpenetration more accurately, we use a new tracer $\tilde{n}$ based on the labeling scheme described in section 2. For each particle species, it is defined as the absolute change in particle density on either side of the shear:
@@ 3 Plasma Mixing and Magnetic Reconnection
To quantify plasma mixing, we compute the volume-averaged $\tilde{n}/n_{0}$ within the $100\times 150\;d_{i}^{2}$ regions highlighted in Figure 2. The left panel of Figure 3 shows its temporal evolution in both shear layers during the nonlinear phase. Although mixing can locally reach $30\%$, its strong localization keeps the volume-averaged fraction low. Ions cross the interface more efficiently than electrons, reaching averages of $\sim 7\%$ compared to $\sim 3\%$ for electrons. Around $t\simeq 300\,\Omega_{c,i}^{-1}$, when fully rolled-up vortices form, mixing increases for both species, especially ions, then plateaus after $t\simeq 850\,\Omega_{c,i}^{-1}$. As Figure 2 shows, low average mixing is expected, since mixing mostly occurs near the interface and within plasma parcels crossing it coherently. We also show here that mixing is similar for different layers, implying that anti–alignment between vorticity and magnetic field plays no evident role in this dynamics.
@@
Figure 3: Temporal evolution of plasma mixing, current density, and X-points. Bottom and top shear layers ($y_{sh,1}$, solid; $y_{sh,2}$, dashed) are indicated in all panels. Left panel: percentage of mixed plasma $\tilde{n}/n_{0}$ for electrons (blue) and ions (red). Middle panel: maximum absolute out-of-plane current density $|J_{z}/(n_{0}q_{i}c)|$ for electrons (blue) and ions (red). Right panel: number of X-points in each shear layer.
@@
To understand the link between magnetic-field dynamics and mixing, we analyzed magnetic-reconnection activity by measuring the out-of-plane current density and X-point formation at the shear layers. The middle panel of Figure 3 shows the temporal evolution of the maximum out-of-plane current density $J_{z}$ for each species and shear layer. In our 2D geometry, the reversal of the in-plane magnetic field is sustained by a $J_{z}$ sharply localized at X-points, making it an effective proxy for magnetic reconnection. $J_{z}$ evolves similarly in both shears, with electron currents generally larger in magnitude, as expected from the mass difference between species. For $t\gtrsim 200\ \Omega_{c,i}^{-1}$, the current density begins to increase in both species, reaching saturation around $t\simeq 400\ \Omega_{c,i}^{-1}$. This coincides with the onset of the early nonlinear KHI and the beginning of the vortex-merging stage (left and middle panels of Figure 1), as well as the previously noted rise in the mixing percentage. The right panel of Figure 3 shows the number of X-points over time at each shear layer. X- and O-points are identified as nulls of the magnetic-flux gradient, $\nabla\psi=\mathbf{0}$, and distinguished by the sign of the Hessian determinant, following standard 2D reconnection diagnostics (Servidio et al. 2009; Wan et al. 2013; Olshevsky et al. 2016). The X-point count increases from $t\gtrsim 300\ \Omega_{c,i}^{-1}$, corresponding to the early nonlinear stage of the KHI. These diagnostics indicate that enhanced reconnection activity—higher $\max|J_{z}|$ and more X-points—coincides with increased plasma mixing ($\tilde{n}/n_{0}$). As reconnection breaks and reconnects field lines, particles from opposite sides of the shear are transported along newly connected lines, promoting cross-field mixing in frontier regions and detached plasma parcels.
@@ 4 Conclusions
We investigated the efficiency of plasma mixing driven by the nonlinear development of the magnetized KHI using high-resolution 2D fully kinetic simulations with finite-Larmor-radius effects (Cerri et al. 2013) and diagnostics including particle tracking, a new mixing tracer, and reconnection proxies. They show that, in this setup, mixing between magnetosheath and magnetospheric plasmas is present but relatively inefficient, even after KH vortices form and the nonlinear stage develops. Mixing is highly localized, reaching $\penalty 10000\ 30\%$ only in sparse locations within a narrow interface layer and isolated advected parcels, and is generally absent farther from the shear layers. On average, ions achieve somewhat higher mixing levels than electrons, whereas electron mixing remains weak and largely confined to localized structures. We can explain this behavior by considering that ions are significantly less magnetized than electrons, allowing them to partially cross the shear interface, whereas electrons remain tightly bound to field lines. With a reduced mass ratio ($m_{i}/m_{e}=64$), the actual electron magnetization in physical plasmas would be even higher, and their mixing correspondingly lower, meaning that electron mixing here likely represents an upper bound. Enhanced mixing occurs mainly within plasma parcels that detach and reattach across the interface, suggesting a significant contribution from magnetic reconnection and individual particle dynamics. Regions located between the interface and these parcels exhibit intermittent increase in mixing, coinciding with enhanced magnetic reconnection signatures, such as elevated out-of-plane current and higher X-point counts. This correspondence suggests that reconnection may facilitate localized cross-field transport during the nonlinear development of the KHI.
While these results provide new insights into localized plasma transport during the nonlinear KHI, there are some limitations that should be noted. The simulations are 2D and use a reduced ion-to-electron mass ratio, which may overestimate electron mobility and limit the range of three-dimensional reconnection dynamics. Despite these limitations, this Letter provides clear, fully kinetic evidence that plasma mixing driven by the nonlinear KHI remains intrinsically localized and heavily influenced by in-plane components of the magnetic field. By introducing a new mixing parameter and combining it with particle tracking and reconnection diagnostics, we identify the specific structures—detached plasma parcels and reconnection sites—where cross-field transport is enhanced. Future work could extend these studies to fully three-dimensional setups, explore realistic mass ratios, and have a more detailed investigation of reconnection physics at the particle level, which could further clarify its role in facilitating cross-field transport.

### v2

# Plasma Mixing Driven by the Collisionless Kelvin–Helmholtz Instability

@@ Abstract
Aims. We investigate the spatial localization, species dependence, and physical mechanisms of plasma mixing driven by the nonlinear evolution of the KHI.
Methods. We perform high-resolution two-dimensional Particle-In-Cell simulations using a finite-Larmor-radius shear-flow initial configuration. Plasma mixing is quantified using particle labeling, a complementary density-based mixing tracer, and diagnostics of magnetic reconnection.
Results. Mixing across the shear layer is present but localized, occurring mainly in narrow interface regions and plasma structures. Ions mix more effectively than electrons, which remain largely frozen to field lines. Enhanced mixing spatially and temporally correlates with localized magnetic reconnection within and between KH vortices.
Conclusions. Cross-boundary transport driven by the kinetic KHI remains intrinsically localized and is mediated by vortex advection and magnetic reconnection. Electron mixing is strongly constrained, indicating that kinetic-scale transport across collisionless shear layers remains limited.
@@ 1 Introduction
The Kelvin–Helmholtz instability (KHI) can develop when the velocity shear exceeds the Alfvén speed associated with the magnetic-field component parallel to the shear (Chandrasekhar 1961). At Earth’s low-latitude magnetopause, this condition is frequently met during intervals of strongly northward or southward interplanetary magnetic field (IMF), and signatures of surface waves and rolled-up vortices have been widely observed (Hasegawa et al. 2004; Eriksson et al. 2016; Stawarz et al. 2016; Settino et al. 2024). KH activity has also been reported at other planetary magnetospheres, including Mercury (Sundberg et al. 2012; Aizawa et al. 2020), Mars (Ruhunusiri et al. 2016; Koh et al. 2025), Saturn (Ma et al. 2015; Dialynas 2018), and Jupiter (Ranquist et al. 2019; Montgomery et al. 2023), raising the question of how magnetospheric size influences the multi-scale development of the instability.
The KHI is also ubiquitous in heliospheric and astrophysical shear flows. In the heliosheath and heliotail, global magnetohydrodynamic (MHD) simulations have shown that velocity shear can drive KH and related instabilities that contribute to large-scale turbulence and plasma transport (e.g., Kieokaew et al. 2021; Ma et al. 2025). Similar shear-driven processes are expected to operate in magnetized astrophysical jets and other collisionless boundary layers, where the nonlinear evolution of KH vortices may influence magnetic-field restructuring and energy dissipation (Chow et al. 2023a, b). These large-scale studies highlight the global impact of shear-driven instabilities, but do not resolve the kinetic-scale processes that regulate particle interpenetration within vortices.
After saturation, KH vortices generate strong gradients in velocity, density, and magnetic field down to ion and electron scales, triggering secondary instabilities that broaden the mixing layer and enhance plasma transport across the solar wind–magnetosphere boundary (Nykyri and Otto 2001; Hasegawa et al. 2004; Nakamura et al. 2017a). Kinetic studies and in-situ observations show that this transport is mediated by processes such as reconnection, wave–particle interactions, and vortex-induced reconnection (VIR; e.g., Nakamura et al. 2011), enabling solar-wind plasma entry even under northward IMF conditions (Nykyri et al. 2006; Stawarz et al. 2016; Nakamura et al. 2017b, 2020). While MHD simulations capture the global KH dynamics and reconnection-driven transport (Faganello et al. 2012; Ma et al. 2017; Ferro et al. 2024), they do not resolve the kinetic mechanisms governing particle diffusion and heating. Despite these advances, quantitatively characterizing plasma mixing in fully kinetic KHI remains challenging, particularly in distinguishing genuine plasma interpenetration from large-scale advection (Nakamura et al. 2013; Karimabadi et al. 2013).
Several studies have addressed plasma transport and mixing in KH dynamics using both simulations and observations. In numerical models, distinguishing plasma populations originating on opposite sides of the shear layer has been achieved through population tagging or passive tracers, to quantify boundary-layer broadening and reconnection-driven transport (e.g., Nakamura and Daughton 2014; Ma et al. 2017). In spacecraft observations, where particle origin cannot be directly tracked, mixing parameters have been introduced based on energy-separated magnetospheric and magnetosheath populations (Settino et al. 2022, 2024; Radhakrishnan et al. 2024). These observational diagnostics provide valuable classification tools for KH-driven particle crossing but do not allow direct measurement of true particle interpenetration or disentangle mixing from heating and spectral evolution.
Observational studies have reported mixed magnetosheath and magnetospheric particle populations within KH vortices and boundary layers, together with signatures of reconnection and VIR (e.g., Hasegawa et al. 2004; Stawarz et al. 2016; Nakamura et al. 2017b; Settino et al. 2022; Radhakrishnan et al. 2024). These measurements provide compelling evidence that KH activity facilitates plasma entry. However, in situ observations are inherently local and do not directly quantify the degree of irreversible cross-field interpenetration over global magnetopause scales. Fully kinetic simulations can therefore help clarify the mechanisms and spatial localization of mixing within KH structures. These simulations provide the possibility to label particles according to their initial spatial origin, enabling a direct quantification of cross-boundary transport. Label-based mixing diagnostics have been introduced in previous PIC studies to investigate plasma entry during KH evolution. In particular, Nakamura and Daughton (2014) performed three-dimensional PIC simulations and quantified plasma transport using a mixing fraction based on particle origin. In that configuration, secondary interchange-type instabilities developed and contributed to turbulent plasma transport across the boundary layer.
The present study examines a different kinetic regime, characterized by an in-plane magnetic field and a symmetric-shear configuration. We focus on how mixing develops during the nonlinear evolution of KH vortices under these conditions, quantifying the mixing spatial localization, species dependence, and temporal evolution, and relating it to kinetic reconnection diagnostics. Rather than estimating global transport rates, our goal is to determine where and how genuine particle interpenetration occurs and whether mixing remains localized at kinetic scales.
To this end, we combine particle labeling with a complementary density-based tracer and reconnection diagnostics to quantify the spatial structure of plasma interpenetration during the nonlinear stage of the KHI using fully kinetic Particle-in-Cell (PIC) simulations of an electron–ion plasma. This framework enables a direct comparison between particle-origin mixing metrics and reconnection activity, allowing us to assess whether and to what extent magnetic reconnection is associated with localized cross-field transport. In contrast to previous studies that primarily inferred plasma transport from large-scale topology changes or global mixing fractions, we focus on the spatial localization and species dependence of genuine particle interpenetration at kinetic scales. By combining particle-origin labeling, a complementary density-based tracer, and reconnection diagnostics, we provide a quantitative assessment of how and where magnetic reconnection activity coincides with enhanced cross-field mixing during the nonlinear evolution of the KHI.
@@ 2 Methodology and Simulation Setup
We model a collisionless magnetized plasma with uniform per-species density $n_{0}$ and an initial magnetic field composed of an in-plane component along $x$ and a guide field $B_{z,0}=10B_{x,0}$. All quantities are normalized using ion (mass $m_{i}$, charge $q_{i}$) reference scales: the ion gyro-frequency $\Omega_{c,i}=q_{i}B_{\mathrm{tot,0}}/(m_{i}c)$, where $B_{\mathrm{tot,0}}=\sqrt{B_{x,0}^{2}+B_{z,0}^{2}}$ is the total magnetic field magnitude at initialization; the ion plasma frequency $\omega_{p,i}=\sqrt{4\pi n_{0}q_{i}^{2}/m_{i}}$; the ion inertial length $d_{i}=c/\omega_{p,i}$; and the Alfvén speed $v_{A}=B_{\mathrm{tot,0}}/\sqrt{4\pi n_{0}m_{i}}$. At initialization, the ion cyclotron-to-plasma frequency ratio is $\Omega_{c,i}/\omega_{p,i}=0.05$.
@@
Plasma mixing is tracked by assigning each particle a label $\ell$ based on its initial $y$-position, with $\ell=0$ for particles initialized in $y<y_{sh,1}$ or $y>y_{sh,2}$, and $\ell=1$ for those initialized in $y_{sh,1}<y<y_{sh,2}$. This labeling allows us to distinguish particles originating from different regions and to track their movement across shear layers, providing a new measure of plasma transport and mixing for both ions and electrons that isolates genuine interpenetration from coherent vortex advection and by systematically quantifying its spatial and temporal evolution at kinetic scales.
3 Results: Plasma Mixing and Magnetic Reconnection
In Figure 1, we show snapshots representative of the evolution of the vortices at the lower shear during the nonlinear stage of the KHI. The quantity shown is the electron mixing fraction (Nakamura and Daughton 2014),
$F_{e}=\frac{(n_{e,\ell=1}-n_{e,\ell=0})}{n_{e}},$ (2)
where $n_{e,\ell=1}$ and $n_{e,\ell=0}$ are the electron number densities initially on either side of the shear. Regions with $F_{e}=-1$, $F_{e}=1$, and $|F_{e}|<1$ correspond to unmixed plasma on either side and mixed plasma, respectively. The ion mixing fraction and the upper shear layer for both species behave similarly and are not shown.
The left panel of Figure 1 shows three rolled-up vortices at the early nonlinear stage. As they merge (middle panel), a dynamic mixing layer forms, where plasma parcels can temporarily cross regions before returning to their origin (Figure 1, middle and left panels). Magnetic tension limits permanent cross-field transport: although reconnection allows plasma parcels to detach from their original field lines, the strong in-plane field pulls them back. Consequently, magnetic tension counteracts plasma transport and forces macroscopic parcels to rebound toward their region of origin, limiting their ability to permanently migrate into the opposite region. In Appendix B we further elaborate on this conjecture, showing how the presence of an in-plane magnetic field influences the development of the KHI and prevents plasma from mixing efficiently.
While the mixing fraction $F_{e}$ provides the qualitative picture described above, it does not allow us to adequately quantify actual mixing because it does not capture the movement of small groups of particles potentially crossing between regions. To quantify interpenetration more accurately, we use a new tracer $\tilde{n}$ based on the labeling scheme described in section 2. For each particle species, it is defined as the absolute change in particle density on either side of the shear:
@@ 3 Results: Plasma Mixing and Magnetic Reconnection
To quantify plasma mixing, we compute the volume-averaged $\tilde{n}/n_{0}$ within the $100\times 150\;d_{i}^{2}$ regions highlighted in Figure 2. The top panel of Figure 3 shows its temporal evolution in both shear layers during the nonlinear phase. Although mixing can locally reach $30\%$, its strong localization keeps the volume-averaged fraction low. Ions cross the interface more efficiently than electrons, reaching averages of $\sim 7\%$ compared to $\sim 3\%$ for electrons. Around $t\simeq 300\,\Omega_{c,i}^{-1}$, when fully rolled-up vortices form, mixing increases for both species, especially ions, then plateaus after $t\simeq 850\,\Omega_{c,i}^{-1}$. As Figure 2 shows, low average mixing is expected, since mixing mostly occurs near the interface and within plasma parcels crossing it coherently. We also show here that mixing is similar for different layers, implying that anti–alignment between vorticity and magnetic field plays no evident role in this dynamics.
@@
Figure 3: Temporal evolution of plasma mixing, current density, and X-points. Bottom and top shear layers ($y_{sh,1}$, solid; $y_{sh,2}$, dashed) are indicated in all panels. Top panel: percentage of mixed plasma $\tilde{n}/n_{0}$ for electrons (blue) and ions (red). Middle panel: maximum absolute out-of-plane current density $|J_{z}/(n_{0}q_{i}c)|$ for electrons (blue) and ions (red). Bottom panel: number of X-points in each shear layer.
@@
To understand the link between magnetic-field dynamics and mixing, we analyzed magnetic-reconnection activity by measuring the out-of-plane current density and X-point formation at the shear layers. Magnetic reconnection has previously been identified as a key process in the nonlinear evolution of KH vortices and suggested as a potential driver of plasma transport (e.g., Hasegawa et al. 2004; Nakamura et al. 2011). Here, we directly quantify its association with particle interpenetration using our mixing tracer. The middle panel of Figure 3 shows the temporal evolution of the maximum out-of-plane current density $J_{z}$ for each species and shear layer. In our 2D geometry, the reversal of the in-plane magnetic field is sustained by a $J_{z}$ sharply localized at X-points, making it an effective proxy for magnetic reconnection. $J_{z}$ evolves similarly in both shears, with electron currents generally larger in magnitude, as expected from the mass difference between species. For $t\gtrsim 200\ \Omega_{c,i}^{-1}$, the current density begins to increase in both species, reaching saturation around $t\simeq 400\ \Omega_{c,i}^{-1}$. This coincides with the onset of the early nonlinear KHI and the beginning of the vortex-merging stage (left and middle panels of Figure 1), as well as the previously noted rise in the mixing percentage. The bottom panel of Figure 3 shows the number of X-points over time at each shear layer. X- and O-points are identified as nulls of the magnetic-flux gradient, $\nabla\psi=\mathbf{0}$, and distinguished by the sign of the Hessian determinant, following standard 2D reconnection diagnostics (Servidio et al. 2009; Wan et al. 2013; Olshevsky et al. 2016). The X-point count increases from $t\gtrsim 300\ \Omega_{c,i}^{-1}$, corresponding to the early nonlinear stage of the KHI. These diagnostics indicate that enhanced reconnection activity—higher $\max|J_{z}|$ and more X-points—coincides with increased plasma mixing ($\tilde{n}/n_{0}$). As reconnection breaks and reconnects field lines, particles from opposite sides of the shear are transported along newly connected lines, promoting cross-field mixing in frontier regions and detached plasma parcels. The spatial correspondence between X-point clusters and regions of elevated mixing, together with their synchronized temporal evolution, provides strong evidence that magnetic reconnection facilitates the localized mixing observed during the nonlinear development of the KHI.
@@ 4 Conclusions
We investigated the efficiency of plasma mixing driven by the nonlinear development of the magnetized KHI using high-resolution 2D fully kinetic simulations with finite-Larmor-radius effects (Cerri et al. 2013) and diagnostics including particle labeling, a new mixing tracer, and reconnection proxies. Unlike previously introduced observational mixing parameters (Settino et al. 2022, 2024), which rely on energy-space separation of particle populations, our diagnostics are based on direct origin tracking in a fully kinetic framework. This enables a quantitative assessment of true particle interpenetration and its localization at electron scales, providing a complementary perspective to spacecraft-based mixing measures. Our results show that, in this setup, mixing between magnetosheath and magnetospheric plasmas is present but relatively inefficient, even after KH vortices form and the nonlinear stage develops. Mixing is highly localized, reaching $\penalty 10000\ 30\%$ only in sparse locations within a narrow interface layer and isolated advected parcels, and is generally absent farther from the shear layers. On average, ions achieve somewhat higher mixing levels than electrons, whereas electron mixing remains weak and largely confined to localized structures. We can explain this behavior by considering that ions are significantly less magnetized than electrons, allowing them to partially cross the shear interface, whereas electrons remain tightly bound to field lines. With a reduced mass ratio ($m_{i}/m_{e}=64$), the actual electron magnetization in physical plasmas would be even higher, and their mixing correspondingly lower, meaning that electron mixing here likely represents an upper bound. Enhanced mixing occurs mainly within plasma parcels that detach and reattach across the interface, suggesting a significant contribution from magnetic reconnection and individual particle dynamics. Regions located between the interface and these parcels exhibit intermittent increase in mixing, coinciding with enhanced magnetic reconnection signatures, such as elevated out-of-plane current and higher X-point counts. We find that enhanced mixing preferentially occurs in regions where reconnection signatures are present, suggesting that magnetic reconnection facilitates localized cross-field particle interpenetration during the nonlinear development of the KHI.
While these results provide new insights into localized plasma transport during the nonlinear KHI, several limitations should be noted. The simulations are two-dimensional and use a reduced ion-to-electron mass ratio, which may overestimate electron mobility and restrict the range of fully three-dimensional reconnection dynamics. In addition, the present simulations adopt a symmetric configuration across the shear layer, whereas the Earth’s magnetopause is typically asymmetric in density, temperature, and magnetic-field strength. Asymmetric configurations can trigger additional interchange-type instabilities and enhance turbulent transport (Nakamura and Daughton 2014; Dargent et al. 2019). The symmetric setup considered here isolates the role of velocity shear, magnetic geometry, and magnetic reconnection in regulating cross-field transport. Our results should therefore be interpreted as characterizing plasma mixing in a magnetized, symmetry-controlled regime, providing a reference case against which more realistic asymmetric configurations can be assessed. While strongly asymmetric conditions may enhance turbulent transport and increase the net amount of cross-boundary plasma exchange, the mechanisms identified here—localized mixing near reconnection sites and within advected plasma parcels—are expected to remain key ingredients of kinetic KHI evolution in asymmetric configurations. The finite horizontal extent of the simulation domain limits the ultimate growth of individual vortices due to periodic boundary conditions. In the reference configuration, vortex merging leads to structures approaching the box size in $x$, which prevents further large-scale horizontal expansion. However, we run additional simulations with larger domains confirming that increasing the number of vortices does not qualitatively alter the localized nature of mixing at kinetic scales. Although longer evolution times could increase the cumulative transport, the strong in-plane magnetic field constrains vertical broadening of the mixing layer, and interpenetration remains confined to narrow regions near reconnection sites and detached plasma parcels. These results should therefore be interpreted as characterizing the kinetic regulation of mixing within KH vortices rather than the long-term global evolution of the magnetopause boundary layer.
Despite these limitations, this study provides clear, fully kinetic evidence that plasma mixing driven by the nonlinear KHI can remain intrinsically localized and strongly influenced by in-plane magnetic-field components. By introducing a complementary mixing tracer and combining it with particle labeling and reconnection diagnostics, we identify the specific structures—detached plasma parcels and reconnection sites—where cross-field transport is enhanced. Our results demonstrate that enhanced mixing preferentially coincides with magnetic-reconnection signatures, establishing a quantitative connection between kinetic reconnection activity and localized plasma interpenetration in nonlinear KH dynamics.
Future work will extend these studies to fully three-dimensional setups, explore realistic mass ratios, and conduct a more detailed investigation of reconnection physics at the particle level, which could further clarify its role in facilitating cross-field transport.